OpenStudy (anonymous):

Solve for X y=x-1 x-y^2=9

OpenStudy (anonymous):

So far Ive done this... x-(x-1)^2=9 then x-(x^2-2x+1)

OpenStudy (anonymous):

=9

OpenStudy (anonymous):

then i reduced it to 3x-x^2=8

OpenStudy (dumbcow):

yes now combine like terms -> -x^2+3x-1 = 9 set equal to zero, move to right side, better to have positive leading coefficient -> x^2-3x +10 = 0 use quadratic formula

OpenStudy (anonymous):

okay, but how do i find x from there?

OpenStudy (dumbcow):

i don't think this has real roots, the answer will have imaginary numbers

OpenStudy (anonymous):

so it has no solution

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

okay thanks

OpenStudy (dumbcow):

practice using the quadratic formula so you can do it on your own next time \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

thank you!!